OpenStudy (anonymous):

what is equivalent to: (3+5(radical3) / (4-2(radical3) *please show steps*

OpenStudy (anonymous):

You would have to rationalize the denominator

OpenStudy (anonymous):

ok, continue?

OpenStudy (anonymous):

\[\frac{ 3+5\sqrt{3} }{ 4-2\sqrt{3}}\] So you would multiply the numerator and the denominator by conjugate of \[4 - 2\sqrt{3}\]

OpenStudy (anonymous):

Do you know how to find the conjugate?

OpenStudy (anonymous):

yes i think can you explain more please

OpenStudy (anonymous):

so the conjugate would be \[4+2\sqrt{3}\] multiply both sides by that and you get \[\frac{ (3+5\sqrt{3})(4+2\sqrt{3}) }{ (4-2\sqrt{3})(4+2\sqrt{3}) }\] Do you think you could simplify that more?

OpenStudy (anonymous):

i think so?

OpenStudy (anonymous):

so it would be: \[\frac{ 12+6\sqrt{3}+20\sqrt{3}+10\sqrt{9} }{ 16+8\sqrt{3}-8\sqrt{3}-4\sqrt{9} }\] right?

OpenStudy (anonymous):

yeah. Can you simplify the radicals/cancel anything out?

OpenStudy (anonymous):

then it would be: \[\frac{ 12+26\sqrt{3}+30 }{ 4 }\] right?

OpenStudy (anonymous):

yeah, you need to add 12 and 30 too.

OpenStudy (anonymous):

so it would be: \[\frac{ 42+26\sqrt{3} }{ 4 }\]

OpenStudy (anonymous):

Yeah, you've got it!

OpenStudy (anonymous):

is there a way to simplify it more?

OpenStudy (anonymous):

divide the top and bottom by 2

OpenStudy (anonymous):

can you divide \[26\sqrt{2}\] by 2

OpenStudy (anonymous):

yes, its just\[13\sqrt{3}\]

OpenStudy (anonymous):

so would it be: \[\frac{ 21+13\sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

and that's it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

No Problem :)